OpenStudy (anonymous):

consider the following quadratic equation: 15x^2+6x=0 using the standard form ax^2+bx+c=0 of the quadratic equation above, factor the left hand side equation into two linear factors

OpenStudy (anonymous):

x=0 15x + 6=0

OpenStudy (anonymous):

= 3x (5x + 2) = 0

OpenStudy (anonymous):

x = 0 x = -2/5

OpenStudy (anonymous):

15x^2+6x=0 3x(5x + 2) = 0